OpenStudy (anonymous):

Evan purchases a board game that is regularly priced at $19.95. It is on sale for 40% off. What is the sale price of the board game?

OpenStudy (anonymous):

@iGreen @IrishBoy123 @ItsZack @arnoldcx @Blonde_Gangsta @brittanydosey @cab70 @christianbruce1996 @fani1996 @fallenangelorchid @Kstate @kels200105 @ObeyRealz @oblivion20 @ortiz09

OpenStudy (anonymous):

if you get a 40% discount, you pay \(100\%-40\%=60\%\) of the price to find \(60\%\) of \(19.95\) multiply \[19.95\times .6\]

OpenStudy (anonymous):

Sale price = Regular price * (1-discount%) discount % of 40% is written as 40%/100%=0.4

OpenStudy (anonymous):

I nvr get sales tax

OpenStudy (anonymous):

you could estimate it as \(20\times 06=2\times 6=12\) so about \(\$12\)

OpenStudy (anonymous):

thats rounded though. dont you need exact?

OpenStudy (anonymous):

i never get sales tax either

OpenStudy (anonymous):

I could ask my mom she has worked at a store and can tell you any sales price without doing anything mathematical at all

OpenStudy (anonymous):

ok plzz

OpenStudy (anonymous):

ok give me a minute :)

OpenStudy (anonymous):

k

OpenStudy (anonymous):

$11.97

OpenStudy (anonymous):

thank u soo much

OpenStudy (anonymous):

welcome you just multiply the total amount by the percent to figure out how much you take off from the total cost then you subtract them and you have the sales price

OpenStudy (anonymous):

I need ur help on more please?